OpenStudy (hulkart):

Find x and y int. 3x + 5y=45

OpenStudy (wcrmelissa2001):

First you need to make y the subject 3x+5y=45 5y=45-3x y=(45-3x)/5 y=9-3/5 x

OpenStudy (hulkart):

Ok

OpenStudy (hulkart):

Hurry

OpenStudy (igreen):

To find the 'x' intercept, plug in 0 for 'y' and solve for 'x'. To find the 'y' intercept, plug in 0 for 'x' and solve for 'y'.

OpenStudy (igreen):

x-intercept: 3x + 5(0) = 45 y-intercept: 3(0) + 5y = 45

OpenStudy (hulkart):

-15,0 as x is one f d options

OpenStudy (hulkart):

Pls ive got a lot more

OpenStudy (wcrmelissa2001):

a lot more of what?